Vector a= 2i+2j+3k
Vector b=i+2j+k
Vector c= 3i+j
Are such that a+yb is perpendicular to c find y.
1
Expert's answer
2019-11-01T15:10:56-0400

Find the vector "y\overrightarrow{b}":


"y\overrightarrow{b}=y\overrightarrow{i}+2y\overrightarrow{j}+y\overrightarrow{k}"

Tnen


"\overrightarrow{a}+y\overrightarrow{b}=(2\overrightarrow{i}+2\overrightarrow{j}+3\overrightarrow{k})+(y\overrightarrow{i}+2y\overrightarrow{j}+y\overrightarrow{k})""\overrightarrow{a}+y\overrightarrow{b}=(2+y)\overrightarrow{i}+(2+2y)\overrightarrow{j}+(3+y)\overrightarrow{k}"

If "\overrightarrow{a}+y\overrightarrow{b}" is perpendicular to "\overrightarrow{c}" , then their dot product is equal to zero:


"(2+y)\cdot 3+(2+2y)\cdot 1+(3+y)\cdot 0=0"

Simplify:


"6+3y+2+2y=0""5y+8=0""y=-\frac{8}{5}""y=-1.6"

Answer: -1.6